# Cohomological Calibration and Curvature Constraints on Product Manifolds: A Topological Lower Bound

### v10

@@ Abstract.
We establish a quantitative relationship between mixed de Rham classes and the geometric complexity of metric connections with totally skew torsion on admissible product manifolds: either both factors are compact oriented surfaces, or $\dim M=4$ with a $(3,1)$ splitting and the mixed class has tensor rank $1$. For any cohomologically calibrated connection $\nabla^{C}$ with non‑parallel torsion $T$, where the harmonic projection of $T$ represents a mixed class $\omega]$, we prove
@@
with $\mathcal{K}$ an intrinsically defined obstruction space. The bound is a topological invariant under metric deformations preserving the parallel‑form strata and provides a sharp obstruction to reducible holonomy. Counterexamples show the hypothesis is optimal.
@@ 1. Introduction
Let $M=M_{1}\times M_{2}$ be a compact oriented product manifold endowed with a product metric $g=g_{1}\oplus g_{2}$, inducing at each point an orthogonal splitting $T_{p}M=V_{1}\oplus V_{2}$. We introduce a topological invariant, the mixed tensor rank of a cohomology class $\omega]\in H^{3}(M;\mathbb{R})$, and show that—after subtracting a correction term that captures the part of the class absorbed by parallel forms—it provides a lower bound for the dimension of the off‑diagonal holonomy subspace generated by any cohomologically calibrated metric connection $\nabla^{C}$ with totally skew torsion.
@@
    $\dim M=4$ with a $(3,1)$ splitting and the mixed class has total tensor rank $1$.
For any admissible $M$, any cohomologically calibrated connection $\nabla^{C}$ with non‑parallel torsion satisfies
@@ 2.1. Intrinsic Definition of Mixed Cohomology Subspaces
    $\dim M=4$ with $(\dim M_{1},\dim M_{2})=(3,1)$ or $(1,3)$, and the mixed class $\omega]$ has tensor rank $\operatorname{rank}_{\mathbb{R}}([\omega]_{\text{mixed}})=1$.
@@
Case (B): $(3,1)$ or $(1,3)$ with $\operatorname{rank}_{\mathbb{R}}([\omega]_{\text{mixed}})=1$. Assume $(\dim M_{1},\dim M_{2})=(3,1)$ (the other case is symmetric). Then $b_{1}(M_{2})=1$ and $H^{2}(M_{2})=0$. Let $\gamma$ be a fixed basis of $\mathcal{H}^{1}(M_{2})$. Any $T\in\mathcal{H}^{2}(M_{1})\otimes\mathcal{H}^{1}(M_{2})$ has rank at most one and can be written uniquely as
@@
Uniqueness follows as in Case (A) because $\gamma$ is non–zero. Hence $\mathcal{V}_{2,1}=\operatorname{span}\{\alpha\}\otimes\mathcal{H}^{1}(M_{2})$ depends only on $\omega]$. The component $(1,2)$ is trivial since $H^{2}(M_{2})=0$, so $\mathcal{V}_{1,2}=\{0\}$ is obviously well defined. ∎
@@ 3. Off-Diagonal Curvature and Holonomy Complexity
Definition 3.2 (Off-diagonal holonomy algebra).
Let $(M,g)$ be a Riemannian product manifold with orthogonal splitting $T_{p}M=V_{1}\oplus V_{2}$, and let $\nabla^{C}$ be a metric connection on $TM$ with totally skew-symmetric torsion. Let $\Pi_{\mathrm{off}}:\mathfrak{so}(T_{p}M)\to\mathfrak{so}(T_{p}M)$ be the linear projection onto off-diagonal endomorphisms:
$\Pi_{\mathrm{off}}(A)(Z):=P_{1}A(P_{2}Z)+P_{2}A(P_{1}Z),$
where $P_{i}:T_{p}M\to V_{i}$ are the orthogonal projections.
Let $\mathcal{G}_{p}:=\{\nabla^{k}R^{C}|_{p}(X_{1},\dots,X_{k};Y,Z)\mid k\geq 0,X_{i},Y,Z\in T_{p}M\}$ be the set of Ambrose–Singer curvature generators of $\mathfrak{hol}_{p}(\nabla^{C})$. Define the off-diagonal curvature span
$\mathcal{S}_{p}^{\mathrm{off}}:=\operatorname{Span}_{\mathbb{R}}\{\Pi_{\mathrm{off}}(A)\mid A\in\mathcal{G}_{p}\}\subset\mathfrak{so}(T_{p}M).$
The off-diagonal holonomy algebra is the Lie subalgebra obtained by intersecting $\mathfrak{hol}_{p}(\nabla^{C})$ with the Lie algebra generated by $\mathcal{S}_{p}^{\mathrm{off}}$ in $\mathfrak{so}(T_{p}M)$:
$\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C}):=\bigl\langle\mathcal{S}_{p}^{\mathrm{off}}\bigr\rangle_{\mathrm{Lie}}\cap\mathfrak{hol}_{p}(\nabla^{C})\subset\mathfrak{hol}_{p}(\nabla^{C}).$
Here $\langle\cdot\rangle_{\mathrm{Lie}}$ denotes the Lie algebra generated in $\mathfrak{so}(T_{p}M)$, and the intersection ensures that $\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})$ is canonically a subalgebra of $\mathfrak{hol}_{p}(\nabla^{C})$. Equivalently, it is the smallest Lie subalgebra of $\mathfrak{hol}_{p}(\nabla^{C})$ whose off-diagonal components contain $\mathcal{S}_{p}^{\mathrm{off}}$.
@@ 5. Main Result and Proof
Let $(M_{1},g_{1})$ be a compact oriented Riemannian manifold and let $\mathcal{H}^{2}(M_{1},g_{1})$ denote the space of harmonic $2$-forms. Given a subspace $U\subset\mathcal{H}^{2}(M_{1},g_{1})$ of dimension $r$ and its parallel part $U_{\parallel}:=U\cap\mathcal{P}_{2}(M_{1},g_{1})$ of dimension $s$, choose a basis $\alpha_{1},\dots,\alpha_{r}$ of $U$ such that $\alpha_{1},\dots,\alpha_{s}$ span $U_{\parallel}$. Then there exists a non‑empty open set $\mathcal{U}\subset M_{1}$ on which the covariant derivatives
@@
2. The real‑analytic case. Assume now that $g_{1}$ is real‑analytic. Then every harmonic form is real‑analytic (by elliptic regularity for the Hodge–Laplacian, see [8]), and consequently the covariant derivatives $\nabla^{g_{1}}\alpha_{i}$ are real‑analytic sections of the vector bundle $E:=T^{*}M_{1}\otimes\Lambda^{2}T^{*}M_{1}$.
@@
Define the analytic $2$-form $\eta:=\sum_{i=1}^{r-s}c_{i}\alpha_{s+i}$ on $V$. From (5.1) we have $\nabla^{g_{1}}\eta=\sum_{i=1}^{r-s}(dc_{i})\otimes\alpha_{s+i}+\sum_{i=1}^{r-s}c_{i}\nabla^{g_{1}}\alpha_{s+i}=\sum_{i=1}^{r-s}(dc_{i})\otimes\alpha_{s+i}$.
Let $V\subset M_{1}$ be the non–empty open subset fixed above. Since the mixed cohomology class on $M_{1}\times M_{2}$ is assumed to be non–trivial, its harmonic representative has a non–vanishing component in $H^{2}(M_{1})\otimes H^{1}(M_{2})$. In particular, the projection onto $H^{2}(M_{1})$ is non–zero. Hence the harmonic forms $\alpha_{s+1},\dots,\alpha_{r}$ cannot all vanish identically on $M_{1}$. Let $W\subset M_{1}$ denote the complement of their common zero set. By real–analyticity, $W$ is a non–empty open subset of $M_{1}$.
Restricting to the non‑empty open set $V^{\prime}:=V\cap W$, we have on $V^{\prime}$
$\sum_{i=1}^{r-s}(dc_{i})\otimes\alpha_{s+i}=0.$
Because the $\alpha_{s+i}(p)$ are linearly independent at every $p\in V^{\prime}$, this forces $dc_{i}=0$ on $V^{\prime}$ for each $i$. Thus the functions $c_{i}$ are constant on each connected component of $V^{\prime}$. By analyticity, they extend to constants on the whole connected open set $V$ (if two analytic functions coincide on an open set, they coincide everywhere on the connected component). Consequently, the $c_{i}$ are constant on $V$. Thus there exist constants $c_{1},\dots,c_{r-s}\in\mathbb{R}$, not all zero, such that
$\sum_{i=1}^{r-s}c_{i}\nabla^{g_{1}}\alpha_{s+i}=0\qquad\text{on }V.$
Define the global harmonic $2$-form
$\zeta:=\sum_{i=1}^{r-s}\bar{c}_{i}\alpha_{s+i}\in\mathcal{H}^{2}(M_{1}).$
Then $\nabla^{g_{1}}\zeta=0$ on $V$; i.e., $\zeta$ is parallel on the open set $V$. Because both $\zeta$ and the condition $\nabla^{g_{1}}\zeta=0$ are real-analytic, the unique-continuation principle [1] implies that $\nabla^{g_{1}}\zeta=0$ on the whole connected component of $M_{1}$ containing $V$, hence on all of $M_{1}$. Therefore $\zeta$ is a globally parallel $2$-form. But this contradicts the hypothesis that $\alpha_{s+1},\dots,\alpha_{r}$ are linearly independent modulo $\mathcal{P}_{2}(M_{1},g_{1})$. Indeed, $\zeta$ is a non-trivial linear combination of these forms that belongs to $\mathcal{P}_{2}(M_{1},g_{1})$. Hence our assumption that $\mathcal{U}=\emptyset$ is impossible, and we conclude that $\mathcal{U}$ is a non-empty open subset of $M_{1}$.
@@
Then, for every $p\in M$,
@@
Consider the curvature formula (2.5). The torsion $T$ decomposes as $T=\omega_{h}+\eta$, where $\omega_{h}$ is the harmonic representative of the mixed class $\omega]$ and $\eta=d\alpha+\delta\beta$ collects the exact and coexact parts.
Our aim is to show that the off‑diagonal holonomy algebra contains at least $r^{\sharp}$ linearly independent directions. These directions will be produced by the linear term $\frac{1}{2}(\nabla^{LC}T)_{\mathrm{off}}$ arising from the harmonic part $\omega_{h}$ alone. We must verify that the remaining contributions—the part of the linear term coming from $\eta$ and the quadratic term $Q_{T}$—cannot interfere with this lower bound.
Suppose that on some non‑empty open set $V\subset M$, the derivative $\nabla^{LC}\eta$ cancelled enough off‑diagonal directions from $\nabla^{LC}\omega_{h}$ to reduce the rank of $(\nabla^{LC}T)_{\mathrm{off}}$ below $r^{\sharp}$. This would impose linear relations among the covariant derivatives of the harmonic components $\alpha_{i}\wedge\beta_{i}$ on $V$. Applying the analytic‑approximation technique employed in Lemma 5.1 (see also Remark 5.6), we may work with real‑analytic metrics and forms. The relation then becomes an equality between real‑analytic sections, so by the unique continuation principle for analytic sections (Aronszajn [1]) it extends to the whole connected component containing $V$, hence globally. Since the tensors $\nabla^{LC}(\alpha_{i}\wedge\beta_{i})$ are pointwise linearly independent on the open set $\mathcal{U}$ provided by Lemma 5.1, evaluating the global relation at any point of $\mathcal{U}$ forces all coefficients to vanish, contradicting the existence of a non‑trivial cancellation. Hence, on a (possibly smaller) non‑empty open set, the linear term $(\nabla^{LC}T)_{\mathrm{off}}$ still exhibits at least $r^{\sharp}$ independent off‑diagonal directions originating from $\omega_{h}$. Regarding the quadratic term $Q_{T}$, observe that the $r^{\sharp}$ independent operators produced by the linear term belong to the off‑diagonal holonomy algebra by the Ambrose–Singer theorem [2]. The quadratic term contributes additional operators to the same algebra. Since the dimension of a vector space generated by a set of vectors is at least the rank of any linearly independent subset, adding further generators cannot reduce the dimension below $r^{\sharp}$. Consequently, the lower bound obtained from the linear term persists for the full curvature $R^{C}$. Thus, to establish $\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}$ it suffices to analyse the linear term generated by the harmonic component $\omega_{h}$.
@@
linearly independent curvature operators. By the Ambrose–Singer theorem [2], these operators belong to the holonomy algebra at $p$. By Definition 3.2 of the off‑diagonal holonomy algebra $\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})$ as the Lie subalgebra generated by all off‑diagonal curvature components and their covariant derivatives, we therefore have
@@
Now let $p_{0}$ be any point in $\mathcal{U}$. Because the manifold $M$ is connected, the holonomy algebras at different points are isomorphic via parallel transport along curves. In particular, the dimension of the off‑diagonal holonomy algebra is constant on $M$. Hence the inequality $\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}$, which holds at $p_{0}$, must hold at every point $p\in M$ and no density requirement is needed. This completes the proof of the lower bound.
The lower bound depends only on $\omega]$ and on the spaces of $g_{1}$‑parallel harmonic 22‑forms and 11‑forms, and of $g_{2}$‑parallel harmonic 11‑forms and 22‑forms; it is therefore invariant under any deformation of $g_{1}$ or $g_{2}$ that preserves these parallel‑form strata, hence under any product‑metric deformation that leaves $\mathcal{P}_{k}(M_{i})$ unchanged for $k=1,2$. ∎
@@
In Theorem 5.2 the analytic approximation is used only indirectly, through Lemma 5.1. Once the lemma guarantees a non‑empty open set where the linear maps $\Psi_{p}$ and $\widetilde{\Psi}_{p}$ have maximal rank, the inequality $\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}$ follows first on that open set, and then on the whole manifold because the dimension of the holonomy algebra is constant on connected components (Ambrose–Singer [2]).

### v11

@@ Abstract.
We establish a quantitative relationship between mixed de Rham classes and the geometric complexity of metric connections with totally skew torsion on admissible product manifolds: either both factors are compact oriented surfaces, or $\dim M=4$ with a $(3,1)$ splitting. For any cohomologically calibrated connection $\nabla^{C}$ with non‑parallel torsion $T$, where the harmonic projection of $T$ represents a mixed class $\omega]$, we prove that on a non-empty open subset $\mathcal{U}\subset M$,
@@
with $\mathcal{K}$ an intrinsically defined obstruction space. The bound is a topological invariant under metric deformations preserving the parallel‑form strata and provides an obstruction to reducible holonomy. Counterexamples show the hypothesis is optimal.
@@ 1. Introduction
Let $M=M_{1}\times M_{2}$ be a compact oriented product manifold [^1] As is customary in Riemannian geometry, all manifolds are assumed to be connected unless otherwise stated. endowed with a product metric $g=g_{1}\oplus g_{2}$, inducing at each point an orthogonal splitting $T_{p}M=V_{1}\oplus V_{2}$. We introduce a topological invariant, the mixed tensor rank of a cohomology class $\omega]\in H^{3}(M;\mathbb{R})$, and show that—after subtracting a correction term that captures the part of the class absorbed by parallel forms—it provides a lower bound for the dimension of the off‑diagonal holonomy subspace generated by any cohomologically calibrated metric connection $\nabla^{C}$ with totally skew torsion.
@@
    $\dim M=4$ with a $(3,1)$ splitting.
For any admissible $M$, any cohomologically calibrated connection $\nabla^{C}$ with non‑parallel torsion satisfies, on a non-empty open subset $\mathcal{U}\subset M$,
@@ 2.1. Intrinsic Definition of Mixed Cohomology Subspaces
    $\dim M=4$ with $(\dim M_{1},\dim M_{2})=(3,1)$ or $(1,3)$, and the mixed class $\omega]$ has tensor rank $\operatorname{rank}_{\mathbb{R}}([\omega]_{\text{mixed}})=1$. [^2] This is not an additional condition, but holds automatically: $\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})\geq 1$ because $\omega]_{\mathrm{mixed}}\neq 0$ (not trivial mixed class), while $\operatorname{rank}_{\mathbb{R}}([\omega]_{\mathrm{mixed}})\leq 1$ since $S^{1}$ is the unique compact connected oriented $1$-manifold, with $\dim H^{1}(S^{1})=1$ and $H^{2}(S^{1})=0$; hence equality holds.
@@
Case (B): $(3,1)$ or $(1,3)$. Assume $(\dim M_{1},\dim M_{2})=(3,1)$ (the other case is symmetric). Then $b_{1}(M_{2})=1$ and $H^{2}(M_{2})=0$. Let $\gamma$ be a fixed basis of $\mathcal{H}^{1}(M_{2})$. Any $T\in\mathcal{H}^{2}(M_{1})\otimes\mathcal{H}^{1}(M_{2})$ has rank at most one and can be written uniquely as
@@
Uniqueness follows as in Case (A) because $\gamma$ is non–zero. Hence $\mathcal{V}_{2,1}=\operatorname{span}\{\alpha\}\otimes\mathcal{H}^{1}(M_{2})$ is intrinsically defined. The component $(1,2)$ is trivial since $H^{2}(M_{2})=0$, so $\mathcal{V}_{1,2}=\{0\}$ is obviously well defined. ∎
@@ 2.2. Mixed obstruction spaces
Remark 2.6 (Anticipatory notation).
The notation $\ker\Psi_{p}$” is anticipatory: in the proof of Theorem 5.2, we construct obstruction maps $\Psi_{p}:V_{2,1}\to\mathfrak{so}(T_{p}M)$ and $\widetilde{\Psi}_{p}:V_{1,2}\to\mathfrak{so}(T_{p}M)$ via covariant derivatives of the torsion form. A key step in the proof is to verify that the kernels of these maps coincide with the intersections defined above. The induced quotient maps $V_{2,1}/\ker\Psi_{p}\to\mathfrak{so}(T_{p}M)$ are then shown to be injective on a non-empty open subset, establishing the lower bound.
@@ 3. Off-Diagonal Curvature and Holonomy Complexity
Definition 3.2 (Off-diagonal holonomy component).
Let $(M,g)$ be a Riemannian product manifold with orthogonal splitting $T_{p}M=V_{1}\oplus V_{2}$, and let $\nabla^{C}$ be a metric connection on $TM$ with totally skew-symmetric torsion.
Let $P_{i}:T_{p}M\to V_{i}$ be the orthogonal projections and let $\Pi_{\mathrm{off}}:\mathfrak{so}(T_{p}M)\to\mathfrak{so}(T_{p}M)$ be the linear projection onto off-diagonal endomorphisms:
$\Pi_{\mathrm{off}}(A)(Z):=P_{1}A(P_{2}Z)+P_{2}A(P_{1}Z).$
Define the space of off-diagonal skew-symmetric endomorphisms as the image of this projection:
$\mathfrak{so}^{\mathrm{off}}(T_{p}M):=\operatorname{Im}(\Pi_{\mathrm{off}})=\{A\in\mathfrak{so}(T_{p}M)\mid A(V_{1})\subset V_{2},\,A(V_{2})\subset V_{1}\}.$
The off-diagonal holonomy component is the intersection of the holonomy algebra with this space:
$\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C}):=\mathfrak{hol}_{p}(\nabla^{C})\cap\mathfrak{so}^{\mathrm{off}}(T_{p}M).$
Equivalently, $\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})=\{A\in\mathfrak{hol}_{p}(\nabla^{C})\mid\Pi_{\mathrm{off}}(A)=A\}$ consists of those holonomy elements that are purely off-diagonal.
Remark 3.3.
Although $\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})$ is a vector subspace of $\mathfrak{hol}_{p}(\nabla^{C})$, it is not closed under the Lie bracket, since $\mathfrak{so}^{\mathrm{off}},\mathfrak{so}^{\mathrm{off}}]\subset\mathfrak{so}(V_{1})\oplus\mathfrak{so}(V_{2})$. Nevertheless, $\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})>0$ implies that the holonomy representation is irreducible with respect to the splitting $V_{1}\oplus V_{2}$.
@@ 5. Main Result and Proof
Let $(M_{1},g_{1})$ be a compact oriented Riemannian manifold and let $\mathcal{H}^{2}(M_{1},g_{1})$ denote the space of harmonic $2$-forms. Given a subspace $U\subset\mathcal{H}^{2}(M_{1},g_{1})$ of dimension $r$ and its parallel part $U_{\parallel}:=U\cap\mathcal{P}_{2}(M_{1},g_{1})$ of dimension $s$, choose a basis $\alpha_{1},\dots,\alpha_{r}$ of $U$ such that $\alpha_{1},\dots,\alpha_{s}$ span $U_{\parallel}$.
If $r-s\leq 1$ (condition satisfied by admissible product manifolds: $U$ is the span of a minimal harmonic decomposition, so $\dim U=r_{2,1}\leq 1$, hence $r-s\leq 1$; see Definition 4.1, Definition 2.2 and Lemma 2.3), then there exists a non‑empty open set $\mathcal{U}\subset M_{1}$ on which the covariant derivatives
@@
2. Admissible manifolds. Assume now that $g_{1}$ is real‑analytic. Then every harmonic form is real‑analytic (by elliptic regularity for the Hodge–Laplacian, see [8]), and consequently the covariant derivatives $\nabla^{g_{1}}\alpha_{i}$ are real‑analytic sections of the vector bundle $E:=T^{*}M_{1}\otimes\Lambda^{2}T^{*}M_{1}$.
@@
Since we have to consider admissible manifolds (Definition 2.2 and Lemma 2.3) we have two cases:
Case $r=s$: There are no non‑parallel forms to consider; the lemma holds vacuously.
Case $r=s+1$: There exists a unique non‑parallel form $\alpha$ and a single analytic function $c$ such that (5.1) reads $c(p)(\nabla^{g_{1}}\alpha)_{p}=0$ for all $p\in V$. If $c(p_{0})\neq 0$ for some $p_{0}\in V$, then by analyticity $c(p)\neq 0$ on an open subset $V^{\prime}\subset V$, whence $(\nabla^{g_{1}}\alpha)_{p}=0$ on $V^{\prime}$. By the unique‑continuation principle for analytic sections (Aronszajn [1]; see also [8]), $\nabla^{g_{1}}\alpha\equiv 0$ on the whole connected component containing $V^{\prime}$, and hence on all of $M_{1}$. This contradicts the assumption that $\alpha$ is non‑parallel modulo $\mathcal{P}_{2}(M_{1},g_{1})$. Therefore $c\equiv 0$ on $V$, contradicting the choice of the $c_{i}$’s. Hence $\mathcal{U}\neq\emptyset$.
@@
The proof uses only properties of harmonic forms that hold for any degree $k$: analytic approximation, semicontinuity of the dimension of parallel forms, $C^{1}$ convergence, and unique continuation. Therefore, the same result holds for $\mathcal{H}^{1}(M_{1},g_{1})$, as well as for $\mathcal{H}^{1}(M_{2},g_{2})$ and $\mathcal{H}^{2}(M_{2},g_{2})$.
@@
Then, there exists a non-empty open subset $\mathcal{U}\subset M$ such that for every $p\in\mathcal{U}$,
@@
Consider the curvature formula (2.5):
$R^{C}=R^{LC}+\frac{1}{2}(\nabla^{LC}T)+\frac{1}{4}Q_{T}.$
The torsion $T$ decomposes as $T=\omega_{h}+\eta$, where $\omega_{h}$ is the harmonic representative of the mixed class $\omega]$ and $\eta=d\alpha+\delta\beta$ collects the exact and coexact parts. Taking off-diagonal components:
$R^{C}_{\mathrm{off}}=R^{LC}_{\mathrm{off}}+\frac{1}{2}(\nabla^{LC}\omega_{h})_{\mathrm{off}}+\frac{1}{2}(\nabla^{LC}\eta)_{\mathrm{off}}+\frac{1}{4}(Q_{T})_{\mathrm{off}}.$
Our aim is to show that $\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}:=r-\dim\mathcal{K}$ on a non-empty open subset of $M$. We first establish that to prove this bound, it suffices to construct $r^{\sharp}$ linearly independent off-diagonal operators from the harmonic term $\frac{1}{2}(\nabla^{LC}\omega_{h})_{\mathrm{off}}$ alone.
Suppose that at some point $p\in M$ we have constructed operators $A_{1},\ldots,A_{r^{\sharp}}\in\mathfrak{so}^{\mathrm{off}}(T_{p}M)$ that are linearly independent and obtained from evaluations of $\frac{1}{2}(\nabla^{LC}\omega_{h})_{\mathrm{off}}$. By the curvature formula, each such operator is a value of $R^{C}_{\mathrm{off}}$, since $(\nabla^{LC}\omega_{h})_{\mathrm{off}}$ appears as one of the terms in $R^{C}_{\mathrm{off}}$. Let
$S_{p}^{\mathrm{off}}:=\mathrm{Span}_{\mathbb{R}}\{\Pi_{\mathrm{off}}(R^{C}(X,Y)):X,Y\in T_{p}M\}$
be the span of all off-diagonal curvature values at $p$. By the curvature formula (2.5), the covariant derivative $(\nabla^{LC}_{X}T)(Y,Z,\cdot)$ appears in the operator $R^{C}(X,Y)$ after raising indices; hence the operators $A_{1},\ldots,A_{r^{\sharp}}$ constructed from evaluations of $(\nabla^{LC}\omega_{h})_{\mathrm{off}}$ are indeed components of curvature values, and therefore belong to $S_{p}^{\mathrm{off}}$. Since they are linearly independent, we have $\dim S_{p}^{\mathrm{off}}\geq r^{\sharp}$. The terms $R^{LC}_{\mathrm{off}}$, $(\nabla^{LC}\eta)_{\mathrm{off}}$, and $(Q_{T})_{\mathrm{off}}$ contribute additional values to $S_{p}^{\mathrm{off}}$, but cannot reduce its dimension below $r^{\sharp}$. By the Ambrose–Singer theorem [2], $\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})$ contains $S_{p}^{\mathrm{off}}$ as the off-diagonal part of the algebra generated by curvature values, whence $\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}$. Thus, to establish the desired lower bound, it suffices to construct $r^{\sharp}$ linearly independent off-diagonal operators from the harmonic component $\omega_{h}$.
@@
(Note that the coefficients $\{c_{i}\}$ are the same in both conditions, being the components of the fixed element $\xi=\sum c_{i}\alpha_{i}\otimes\beta_{i}$ in the tensor product decomposition.)
@@
linearly independent curvature operators. By the Ambrose–Singer theorem [2], the holonomy algebra $\mathfrak{hol}_{p}(\nabla^{C})$ is generated by curvature values. The off-diagonal component $\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})=\mathfrak{hol}_{p}(\nabla^{C})\cap\mathfrak{so}^{\mathrm{off}}(T_{p}M)$ (Definition 3.2) contains all off-diagonal projections of elements of $\mathfrak{hol}_{p}(\nabla^{C})$. Since the $r^{\sharp}$ operators constructed above are off-diagonal curvature operators, they belong to $\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})$, and therefore
@@
In particular, if $r^{\sharp}>0$, the holonomy representation is irreducible with respect to $V_{1}\oplus V_{2}$.
The lower bound depends only on $\omega]_{\mathrm{mixed}}$ and on the parallel-form strata $\mathcal{P}_{k}(M_{i})$; it is therefore invariant under any product-metric deformation that preserves these strata. ∎
@@
In Theorem 5.2 the analytic approximation is used only indirectly, through Lemma 5.1. Once the lemma guarantees a non‑empty open set where the linear maps $\Psi_{p}$ and $\widetilde{\Psi}_{p}$ have maximal rank, the inequality $\dim\mathfrak{hol}_{p}^{\mathrm{off}}(\nabla^{C})\geq r^{\sharp}$ follows on that open set.